# DATABench: Evaluating Dataset Auditing in Deep Learning from an Adversarial Perspective

### v2

@@ Abstract
The widespread application of Deep Learning across diverse domains hinges critically on the quality and composition of training datasets. However, the common lack of disclosure regarding their usage raises significant privacy and copyright concerns. Dataset auditing, which aims to determine if a specific dataset was used to train a given suspicious model, provides promising solutions to addressing these transparency gaps. While prior work has developed various auditing methods, their resilience against dedicated adversarial attacks remains largely unexplored. To bridge the gap, this paper initiates a comprehensive study evaluating dataset auditing from an adversarial perspective. We start with introducing a novel taxonomy, classifying existing methods based on their reliance on internal features (IF) (inherent to the data) versus external features (EF) (artificially introduced for auditing). Subsequently, we formulate two attack types: evasion attacks, designed to conceal the use of a dataset, and forgery attacks, intending to falsely implicate an unused dataset. Building on the understanding of existing methods and attack objectives, we further propose systematic attack strategies: decoupling, removal, and detection for evasion; adversarial example-based methods for forgery. These formulations and strategies lead to our new benchmark, DATABench, comprising 17 evasion attacks, 5 forgery attacks, and 9 representative auditing methods. Extensive evaluations using DATABench reveal that none of the evaluated auditing methods are sufficiently robust or distinctive under adversarial settings. These findings underscore the urgent need for developing a more secure and reliable dataset auditing method capable of withstanding sophisticated adversarial manipulation.
@@ 1 Introduction
The rapid advancement of Deep Learning (DL) has propelled it into widespread applications across various real-world systems, such as computer vision [34], artificial intelligence (AI) chatbots [30], and self-driving vehicles [11]. Currently, DL models primarily make predictions based on their knowledge learned from a dataset. As such, its success highly relies on and is inseparable from high-quality datasets [68, 46]. As a critical resource and fundamental component of DL, datasets largely determine the performance of the models trained on them. However, developers of DL models and applications usually do not disclose the training datasets they use. This lack of transparency gives rise to potential privacy and copyright concerns. For instance, training data may include unauthorized private information. Commercially deployed models may have been trained using copyrighted datasets that prohibit commercial use. Additionally, many existing data protection regulations, such as GDPR, have granted data owners the right to be informed of their data usage. Consequently, auditing the datasets used in DL models has become a critical issue for safeguarding data privacy and copyright [42, 20].
Dataset auditing [^1] Existing literature may use different terminologies, such as dataset ownership verification [51], dataset copyright auditing [20], or data-use auditing [42]. In this paper, we standardize the term as dataset auditing in DL. in DL [51, 20, 43] refers to determining whether a specific dataset is used to train a suspicious model. Typically, as shown in Figure 1, there are three stages involved in dataset auditing [54]. In the first stage, the data owner collects some data from various resources. Subsequently, they process the dataset ($e.g.$, data annotation) to make it suitable for usage or release. Then, in the second stage, an adversary may get access to the dataset through copying or stealing and utilize it to train a DL model without authorization. Third, if a third-party DL model (dubbed a ‘suspicious model’) is suspected of using a protected dataset unauthorizedly, dataset auditing techniques will be adopted to verify it. However, there exist several challenges to achieving a reliable and accurate dataset auditing. (1) It is hard to uncover the traces of a dataset only from the parameters, intermediate results, or even black-box access to a trained DL model. These revealed clues may not be directly associated with the data. (2) The adversary can use various different model architectures and training techniques to produce the model. The diversity and the training-agnostic situation make it difficult to achieve robust and distinctive dataset auditing. (3) The adversary may intentionally process the data or model to compromise the auditing procedure, challenging the robustness of data auditing solutions. As such, achieving dataset auditing is a non-trivial issue in practice [20, 71, 14].
Although some existing works have proposed promising solutions to dataset auditing [79, 31, 61], there is still a significant gap between how the robustness and distinctiveness of existing works are evaluated (often in controlled environments) and how they perform in real-world practice. Existing dataset auditing research primarily considers naive operations and techniques in DL under constrained settings, such as training with models of different architectures, fine-tuning, and pruning, as potential adversarial vectors [51, 20, 33]. These works lack the consideration of a ‘clever’ adversary who may design specified and adaptive attacks against dataset auditing. It is challenging to conduct an exhaustive evaluation due to the complexity of training a model in DL and the diversity of attack surfaces. Consequently, the following critical question remains unanswered: are existing dataset auditing methods reliable under adversarial settings in practice?
@@
  • Contribution 2: Unified Formulations of Potential Attacks against Dataset Auditing.
@@
2 Dataset Auditing in Deep Learning
2.1 Preliminaries
Deep Learning. Deep Learning (DL) has become the most famous and widely-used technique to develop an artificial intelligence (AI) application during the last decade [34, 68], especially with the growth of large foundation models [30]. DL primarily leverages deep neural networks (DNNs) as the core model architecture to achieve a wide range of functions.
While developing a DL model $\mathcal{M}$, the model developer needs to first collect a dataset $\mathcal{D}=\{(\bm{x}_{i},\bm{y}_{i})\}_{i=1}^{|\mathcal{D}|}$. $\mathcal{D}$ is related to the developer-specified DL task and consists of various input samples $\bm{x}_{i}$ and their corresponding label $\bm{y}_{i}$. Subsequently, the developer designs a loss function $\mathcal{L}(\cdot)$ and optimizes the model $\mathcal{M}$ according to the following optimization problem:
@@ 2.1 Preliminaries
Privacy and Copyright Concern about DL Datasets. Currently, most of the DL techniques and models are data-centric, and their developments highly rely on a high-quality dataset [68]. This reliance, however, raises significant concerns regarding data privacy and copyright. For instance, many DL models are trained on large-scale datasets that may contain sensitive or proprietary information, leading to potential violations of user privacy and intellectual property rights [7, 20]. However, classical data protection techniques, such as encryption [65], differential privacy [1], traditional data watermarking [87], and unlearnable data [40], are not sufficient for resolving this issue (details in Appendix J). Consequently, dataset auditing, as a technique that has negligible influence on the data utility and can protect the privacy and copyright in DL scenarios, has become an emerging yet important solution to data protection [48, 20].
2.2 Towards a Formal Definition of Dataset Auditing in DL
@@
Before defining dataset auditing, we hereby first introduce the general threat model of dataset auditing in DL [51, 20]. Specifically, there are two parties involved in the threat model, the data owner and the adversary [^2] Some existing literature [57, 76, 20] also introduces a trustworthy third-party auditor (or verifier) in their threat model to conduct the audit. However, since the auditor does not affect the technical design of dataset auditing methods, we exclude it to avoid potential confusion.. The data owner wants to collect, store, and publish a DL dataset. However, there exists the risk that an adversary may unlawfully copy or steal the dataset and train their own model for unauthorized usage. Such a misbehavior may compromise the privacy or intellectual property rights of the data owner. As such, the data owner seeks an effective dataset auditing method to confirm whether the suspicious model of the adversary is trained on their dataset. We formally define dataset auditing as follows.
@@
The auxiliary information $\mathcal{I}$ in Definition 1 varies depending on the specific dataset auditing method. Common forms of $\mathcal{I}$ include auxiliary datasets used for auditing [77], specific trigger patterns [51], and watermark information [42].
Notably, if the number of samples in $\mathcal{D}$ is 11 ($i.e.$, $|\mathcal{D}|=1$), the definition of dataset auditing is similar to membership inference [7, 66, 74]. As such, existing membership inference techniques can be easily adapted to achieve dataset auditing, which will be further discussed in Section 2.3.
@@
According to the operation and capability of the data owner, existing dataset auditing methods can be further classified according to the following simple taxonomies [20].
@@
In the preparation stage, if the data owner does not need to alter the raw dataset, $i.e.$, $\mathcal{D}=\widehat{\mathcal{D}}$, the dataset auditing method is a non-intrusive method. Otherwise, it is an intrusive one.
@@
In the auditing stage, if $\mathcal{A}$ only needs API access to the suspicious model ($i.e.$, only get the outputs of given inputs), it is defined as a black-box method. Otherwise, if additional information about the suspicious model is needed ($e.g.$, the parameters or the intermediate results during inference), the dataset auditing method $\mathcal{A}$ is a white-box method.
@@ 2.3 Dataset Auditing Taxonomy
Existing works [20] primarily introduce simple taxonomies such as Definitions 2 or 3. However, these taxonomies offer limited aid in understanding the underlying insights of dataset auditing methods and designing effective attack strategies to evaluate their reliability. As such, we propose a new taxonomy based on the origin of the features exploited during the auditing process. Specifically, we split the features of the data into internal and external features and classify existing dataset auditing methods into the following two main categories.
@@
Internal Feature (IF)-based Dataset Auditing. This class of methods relies on the internal feature of the dataset. Internal features refer to features that naturally exist within the data itself. Learning from these features builds up the utility of the model on the primitive task. The core insight of IF-based dataset auditing is to exploit the characteristics of the internal features ($e.g.$, overfitting) during or after training.
@@
Dataset auditing methods based on membership inference attacks (MIA) [77, 36, 56] fall into this category. MIA aims to determine whether a given data instance was included in a model’s training dataset. Given a trained model and a specific data point, MIA outputs a confidence score that reflects the likelihood that the instance was part of the training set. Although originally proposed as a privacy attack, MIA can be readily adapted to the task of dataset auditing by aggregating the confidence scores across all the samples in a dataset ($e.g.$, calculating the average confidence score).
@@
Additionally, some existing works explore auxiliary mechanisms to enhance the effectiveness of MIA in the context of dataset auditing. Some studies propose Dataset Inference (DI) [61, 25], which treats the dataset as a whole and utilizes the prediction margin to conduct dataset auditing. [60] also explores aggregating results of different MIAs. Data-use Auditing (DUA) [42] adds perturbations to craft two datasets that are significantly different from each other in the feature space. DUA achieves dataset auditing by testing whether the confidence of being a member on one dataset is significantly larger than the other. These mechanisms make the IF-based dataset auditing more feasible in practice.
@@
External Feature (EF)-based Dataset Auditing. In contrast, EF-based dataset auditing methods introduce external or artificial features into the dataset. The model trained on the protected dataset with external features will have identifiable behavior that is hardly learned from other datasets. This class of methods achieves dataset auditing by verifying whether the suspicious model can trigger the designated behavior.
@@
Existing methods have explored different ‘identifiable behaviors’ for auditing. Radioactive Data [71, 3] induces distributional shifts in feature space. Misclassification on specific samples ($i.e.$, backdoor behaviors) is also widely used [20]. A line of research utilizes backdoor techniques [73, 54, 51, 79, 31] to mark the dataset. Any model trained on the marked dataset will misclassify samples containing specific trigger patterns. Instead, Domain Watermark (DW) [33] introduces EFs that can make the model correctly classify some ‘hard’ samples with high accuracy. DW does not intentionally lead to misclassification and thus is more harmless.

### v3

@@ Abstract
The widespread application of Deep Learning across diverse domains hinges critically on the quality and composition of training datasets. However, the common lack of disclosure regarding their usage raises significant privacy and copyright concerns. Dataset auditing, which aims to determine if a specific dataset was used to train a given suspicious model, provides promising solutions to addressing these transparency gaps. While prior work has developed various auditing methods, their resilience against dedicated adversarial attacks remains largely unexplored. To bridge the gap, this paper initiates a comprehensive study evaluating dataset auditing from an adversarial perspective. We start with introducing a novel taxonomy, classifying existing methods based on their reliance on internal features (IF) (inherent to the data) versus external features (EF) (artificially introduced for auditing). Subsequently, we formulate two attack types: evasion attacks, designed to conceal the use of a dataset, and forgery attacks, intending to falsely implicate an unused dataset. Building on the understanding of existing methods and attack objectives, we further propose systematic attack strategies: decoupling, removal, and detection for evasion; adversarial example-based methods for forgery. These formulations and strategies lead to our new benchmark, DATABench, comprising 17 evasion attacks, 5 forgery attacks, and 9 representative auditing methods. Extensive evaluations using DATABench reveal that none of the evaluated auditing methods are sufficiently robust or distinctive under adversarial settings. These findings underscore the urgent need for developing a more secure and reliable dataset auditing method capable of withstanding sophisticated adversarial manipulation. Code is available in https://github.com/shaoshuo-ss/DATABench.
keywords:
Dataset Auditing, Copyright Protection, Privacy, Evasion Attack, Trustworthy AI
@@ 1 Introduction
The rapid advancement of Deep Learning (DL) has propelled it into widespread applications across various real-world systems, such as computer vision [he2016deep], artificial intelligence chatbots [guo2025deepseek], and self-driving vehicles [chen2024end]. Currently, DL models primarily make predictions based on their knowledge learned from a dataset. As such, its success highly relies on and is inseparable from high-quality datasets [radford2021learning, krauss2024verify]. As a critical resource and fundamental component of DL, datasets largely determine the performance of the models trained on them. However, developers of DL models and applications usually do not disclose the training datasets they use. This lack of transparency gives rise to potential privacy and copyright concerns. For instance, training data may include unauthorized private information. Commercially deployed models may have been trained using copyrighted datasets that prohibit commercial use. Additionally, many existing data protection regulations, such as GDPR, have granted data owners the right to be informed of their data usage. Consequently, auditing the datasets used in DL models has become a critical issue for safeguarding data privacy and copyright [huang2024general, du2025sok].
Dataset auditing [^1] Existing literature may use different terminologies, such as dataset ownership verification [li2022untargeted], dataset copyright auditing [du2025sok], or data-use auditing [huang2024general]. In this paper, we standardize the term as dataset auditing in DL. in DL [li2022untargeted, du2025sok, huang2025instance] refers to determining whether a specific dataset is used to train a suspicious model. Typically, as shown in Figure 1, there are three stages involved in dataset auditing [li2023black]. In the first stage, the data owner collects some data from various resources. Subsequently, they process the dataset ($e.g.$, data annotation) to make it suitable for usage or release. Then, in the second stage, an adversary may get access to the dataset through copying or stealing and utilize it to train a DL model without authorization. Third, if a third-party DL model (dubbed a ‘suspicious model’) is suspected of using a protected dataset unauthorizedly, dataset auditing techniques will be adopted to verify it. However, there exist several challenges to achieving a reliable and accurate dataset auditing. (1) It is hard to uncover the traces of a dataset only from the parameters, intermediate results, or even black-box access to a trained DL model. These revealed clues may not be directly associated with the data. (2) The adversary can use various different model architectures and training techniques to produce the model. The diversity and the training-agnostic situation make it difficult to achieve robust and distinctive dataset auditing. (3) The adversary could intentionally process the data or model to compromise the auditing procedure, challenging the robustness of data auditing solutions. As such, achieving reliable dataset auditing is a non-trivial issue in practice [du2025sok, sablayrolles2020radioactive, chen2024catch].
@@
Although several existing works have proposed promising solutions to dataset auditing [tang2023did, guo2024zeromark, maini2020dataset], there is still a significant gap between how the robustness and distinctiveness of existing works are evaluated (often in controlled environments) and how they perform in real-world practice. Existing dataset auditing research primarily considers naive operations and techniques in DL under constrained settings, such as training with models of different architectures, fine-tuning, and pruning, as potential adversarial vectors [li2022untargeted, du2025sok, guo2023domain]. These works lack the consideration of a ‘clever’ adversary who may design specified and adaptive attacks against dataset auditing. It is challenging to conduct an exhaustive evaluation due to the complexity of training a model in DL and the diversity of attack surfaces. Consequently, the following critical question remains unanswered: are existing dataset auditing methods reliable under adversarial settings in practice?
@@
  • Contribution 2: Unified Formulations of Attacks against Dataset Auditing.
@@
2 Background
2.1 Deep Learning
Deep Learning (DL) has become the most famous and widely-used technique to develop an artificial intelligence (AI) application during the last decade [he2016deep, radford2021learning], especially with the growth of large foundation models [guo2025deepseek]. DL primarily leverages deep neural networks (DNNs) as the core model architecture to achieve a wide range of functions.
During the development of a DL model $\mathcal{M}$, the model developer needs to first collect a dataset $\mathcal{D}=\{(\bm{x}_{i},\bm{y}_{i})\}_{i=1}^{|\mathcal{D}|}$. The dataset $\mathcal{D}$ is related to the developer-specified DL task. The dataset consists of various input samples $\bm{x}_{i}$ and their corresponding label $\bm{y}_{i}$. Subsequently, the developer designs a loss function $\mathcal{L}(\cdot)$ and optimizes the model $\mathcal{M}$ according to the following optimization problem:
@@ 2.1 Deep Learning
Currently, most of the DL techniques and models are data-centric and their developments highly rely on a high-quality dataset [radford2021learning]. This reliance, however, raises significant concerns regarding data privacy and copyright. For instance, many DL models are trained on large-scale datasets that may contain sensitive or proprietary information, leading to potential violations of user privacy and intellectual property rights [carlini2022membership, du2025sok]. Consequently, data protection and building high-quality datasets are equally important in practice.
2.2 Data Protection
Data protection has long been a critical aspect of information management [meurisch2021data, li2025rethinking, luo2026shadow, zhu2025stealing]. In the last several decades, researchers have developed many different techniques to achieve various requirements of data protection, as follows.
Encryption. Encryption-based techniques aim to transform the sensitive data into unreadable ciphertext, and any unauthorized party without the secret key cannot access the data [bhanot2015review, fiege1987zero]. Advancements in encryption-based techniques like homomorphic encryption (HE) [paillier1999public, chielle2025recurrent] allow computations on encrypted data, making it promising to be adopted in DL systems. However, encryption-based methods suffer from the limitations of excessive latency and resource demands.
Differential Privacy (DP). DP ensures that the inclusion or exclusion of any single data point in a dataset does not significantly affect the output [dwork2006calibrating]. This is typically achieved by injecting controlled noise into data, queries, or model training processes [dwork2006calibrating, abadi2016deep]. While DP provides strong, mathematically provable privacy guarantees, its application in DL introduces significant challenges. For instance, the added noise often degrades model utility. The trade-off between privacy and utility makes DP a powerful but costly solution for privacy-preserving DL [abadi2016deep, yu2019differentially].
Traditional Data Watermarking. Traditional data watermarking techniques help track ownership and usage of data by embedding hidden markers ($i.e.$, watermark) into the data [hussain2022faststamp, zhang2025icmarks]. The data owner can extract the watermark from the data and prove their ownership. Multiple solutions to embedding a watermark into data have been proposed, such as embedding watermarks into the frequency domain or utilizing DL models [cayre2005watermarking, baluja2017hiding]. Recently, some works have also explored using data watermarking to mark AI-generated content and models [zhang2024remark, shao2025explanation, krauss2024clearstamp, ren2024sok]. However, traditional data watermarking techniques may not be feasible for dataset auditing since the watermark in the data can hardly be inherited by the DL models and their owners will not disclose their training details (especially their training data).
Unlearnable Data. This technique aims to add perturbations to the data so that the DL model cannot learn from it during training [huang2021unlearnable, wang2025provably]. The model trained on the unlearnable dataset will have a low utility. However, in the scenario of protecting a public dataset where we intend to prevent unauthorized usage instead of making it completely useless, these techniques may not be feasible.
In general, previous data protection techniques either have a significant impact on the data and model utility ($e.g.$, HE, DP, and unlearnable data) or are not feasible ($e.g.$, traditional data watermarking) in DL scenarios [li2022untargeted, guo2023domain]. Consequently, dataset auditing, as a technique that has negligible influence on the data utility and can protect the privacy and copyright of data owners, has become an emerging yet important and promising solution to data protection [li2025towards, du2025sok].
3 Dataset Auditing in DL
3.1 Towards a Formal Definition of Dataset Auditing
@@
Before defining dataset auditing, we hereby first introduce the general threat model of dataset auditing in DL [li2022untargeted, du2025sok]. Specifically, there are two parties involved in the threat model, the data owner and the adversary [^2] Some existing literature [liu2024false, shao2025fit, du2025sok] also introduces a trustworthy third-party auditor (or verifier) in their threat model to conduct the audit. However, since the auditor does not affect the technical design of dataset auditing methods, we exclude it to avoid potential confusion.. The data owner wants to collect, store, and publish a DL dataset. However, there exists the risk that an adversary may unlawfully copy or steal the dataset and train their own model for unauthorized usage. Such a misbehavior may compromise the privacy or intellectual property rights of the data owner. As such, the data owner seeks an effective dataset auditing method to confirm whether the suspicious model of the adversary is trained on their dataset. We formally define dataset auditing as follows.
@@
The auxiliary information $\mathcal{I}$ in Definition 1 varies depending on the specific dataset auditing method. Common forms of $\mathcal{I}$ include auxiliary datasets used for auditing [shokri2017membership], specific trigger patterns [li2022untargeted], and watermark information [huang2024general].
Notably, if the number of samples in $\mathcal{D}$ is 11 ($i.e.$, $|\mathcal{D}|=1$), the definition of dataset auditing is similar to membership inference [carlini2022membership, peng2025diffence, shang2025defending]. As such, existing membership inference techniques can be easily adapted to achieve dataset auditing, which will be further discussed in Section 3.2.
@@
According to the operation and capability of the data owner, existing dataset auditing methods can be further classified according to the following simple taxonomies [du2025sok].
@@
In the preparation stage, if the data owner does not need to alter the raw dataset, $i.e.$, $\mathcal{D}=\widehat{\mathcal{D}}$, the dataset auditing method is a non-intrusive method, otherwise an intrusive one.
@@
In the auditing stage, if $\mathcal{A}$ only needs API access to the suspicious model ($i.e.$, only get the outputs of given inputs), it is defined as a black-box method. Otherwise, if additional information about the suspicious model is needed ($e.g.$, the parameters or the intermediate results during inference), $\mathcal{A}$ is a white-box method.
@@ 3.2 Dataset Auditing Taxonomy
Existing works [du2025sok] primarily introduce simple taxonomies such as Definitions 2 or 3. However, these taxonomies offer limited aid in understanding the underlying insights of dataset auditing methods and designing effective attack strategies to evaluate their reliability. As such, we propose a new taxonomy based on the origin of the features exploited during the auditing process. Specifically, we split the data features into internal and external features and classify existing dataset auditing methods into the following two main categories.
@@
Internal Feature-based Dataset Auditing. This class of methods relies on the internal feature (IF) of the dataset. IFs refer to features that naturally exist within the data itself. Learning from these features builds up the utility of the model on the primitive task. The core insight of IF-based methods is to exploit the characteristics of the IFs ($e.g.$, overfitting) during or after training.
@@
Dataset auditing methods based on membership inference attacks (MIA) [shokri2017membership, he2024difficulty, liu2022your] fall into this category. MIA aims to determine whether a given data instance was included in a model’s training dataset. Given a trained model and a specific data point, MIA outputs a confidence score that reflects the likelihood that the instance was part of the training set. Although originally proposed as a privacy attack, MIA can be readily adapted to the task of dataset auditing by aggregating the confidence scores across all the samples in a dataset ($e.g.$, calculating the average confidence score).
@@
Additionally, some existing works explore auxiliary mechanisms to enhance the effectiveness of MIA in the context of dataset auditing. Some studies propose Dataset Inference (DI) [maini2020dataset, dziedzic2022dataset], which treats the dataset as a whole and utilizes the prediction margin to conduct dataset auditing. [maini2024llm] also explores aggregating results of different MIAs. Data-use Auditing (DUA) [huang2024general] adds perturbations to craft two datasets that are significantly different from each other in the feature space. DUA achieves dataset auditing by testing whether the confidence of being a member on one dataset is significantly larger than the other. These mechanisms make the IF-based dataset auditing more feasible in practice.
@@
External Feature-based Dataset Auditing. In contrast, external feature (EF)-based methods introduce external or artificial features into the dataset. The model trained on the protected dataset with external features will have identifiable behavior that is hardly learned from other datasets. This class of methods achieves dataset auditing by verifying whether the suspicious model can trigger the designated behavior.
@@
Existing methods have explored different ‘identifiable behaviors’ for auditing. Radioactive Data [sablayrolles2020radioactive, atli2022effectiveness] induces distributional shifts in feature space. Misclassification on specific samples ($i.e.$, backdoor behaviors) is also widely used [du2025sok]. A line of research utilizes backdoor techniques [shan2022poison, li2023black, li2022untargeted, tang2023did, guo2024zeromark] to mark the dataset. Any model trained on a marked dataset will misclassify samples that contain specific trigger patterns. In contrast, Domain Watermark (DW) [guo2023domain] introduces EFs that enable the model to correctly classify certain hard samples with high accuracy. DW does not deliberately induce misclassification and is therefore inherently harmless.
Differences with Prior Taxonomy. The intrusive and non-intrusive taxonomy describes the owner-side operations, whereas our IF/EF taxonomy is defined by the inherent nature of auditing signals. This distinction is critical because it directly shapes the design space and feasibility of attack strategies (details in Section 4), given that IFs cannot be completely removed by adversaries. For example, DUA [huang2024general] highlights the limitations of the prior taxonomy. It does not introduce any identifiable EF signal; instead, it amplifies the discriminability of IFs to enhance membership evidence through perturbation injection. Therefore, although DUA is categorized as intrusive, it remains fundamentally IF-based.